OpenStudy (anonymous):

Suppose you want to build a small walkway around your garden so that there is someplace to walk around and work on your vegetables. The garden measures 20 feet by 30 feet and you want to increase the area of the garden to 1,000 ft2. What is the width of the walkway? A. Approximately 3.3 feet B. Approximately 3.5 feet C. Approximately 3.7 feet D. Approximately 3.9 feet

OpenStudy (anonymous):

Since the garden is 600 sq ft., you need to add 400 sq ft. Let x = width of walkway. Area of the 4 corners = 4 * x^2 Area of the sides: 2 * (30x + 20x) = 100x Total area: 4x^2 + 100x = 400 Solve: x^2 + 25x - 100 = 0 x = about 3.5 feet